From GlueXWiki
Jump to: navigation, search

January 23, 2014 FDC meeting


  1. Installation [1],[2] (all)
  2. Engineering (Bill)
  3. Electronics [3] (Chris, Nick)
  4. Mini-DAQ results FDC E-log (Beni)
  5. Other


Participants: Eugene, Fernando, Dave, Nick, Chris, Simon, Beni, and Lubomir.


- The third quadrant (down left) has been connected to the detector side (see pictures attached). To keep the cables flat on the mesh, Casey uses long zip ties; we want to make sure that at the inner side of the mesh (second picture) these ties do not touch the CDC. Starting Monday, Nick and Chris (with the help of Casey and Tina if needed) will start routing these quadrant below the platform to the electronics side. Before that, over the weekend, Beni will test the second half of the quadrant (packages 3 and 4) with the mini-DAQ.

- The HV was stable at +2100V/-500V without any trips from last Thursday (1/16) till Tuesday (1/21). One HV channel has a current oscillating between 0.5 and 1 uA, indicating some small discharging every 10-15 min, but never trips. On Tuesday Lubomir increased it to +2200V/-500V and it was stable over the night for all the channels, except four channels where the negative HV had to be lowered by 100V to have it stable. On Wednesday morning suddenly all the positive HV channels were turned off and it is not clear how, because the lab was closed and there was nobody in the Hall. We suspect it was a power glitch, but then why the negative HV channels were still on. Beni suggested checking if the ROCs did restart at the same time which may suggest it was a power problem.

- On Wednesday also the gas bottle was empty. Since there were no techs to crane a new bottle, Lubomir switched to another bottle used before which was empty by Thursday morning. It turned out that one bottle (Ar/CO2 40/60 pre-mixed) at 800 sccpm is good for four days. We decided with Tom to have always one extra, full bottle, craned on the platform, so that we can change the gas without the techs being there. Another option would be to mix the gas using a temporary gas system, but Eugene suggested we better use pre-mixed bottles till the final gas system is ready. Then, as Beni explained, the idea is to have 12 Ar bottles used at the same time (connected in one by the supplier). This is not the case for the CO2 bottles, but we will have many of them at the same time.


- Bill was transporting (that's why not at the meeting) the extra cathode mounted on the strong back to be scanned at a Williamsburg company. It turned out they don't have an optical CMM with these sizes, but suggested to contact the same company (New River Kinetics) at Richmond.


- Grounding: Fernando and Bill decided this morning to use the thick ground cable (4/0, green) only up to the face of the magnet, but then inside the magnet to the FDC to use thinner cable (2/0, orange) instead. The lug for the latter cable is small enough so that it can be attached to the ground plate along the packages. Fernando raised the question of grounding the mesh. Now it is attached to the FDC ground by two plates needed for the mechanical connection: one of them should be insulated while the other one should make electrical contact.

- We discussed if the current through the LV cables (~2A per cable) inside the magnet can result in a force that can do some damage like disconnecting LV connectors. Fernando: the return current goes through the same cables, so the force should be zero. Eugene has a different practical experience seing LV cables moving inside the magnet in a CERN experiment.

- Chris: the adapter (a PCB) for the 3M connector at the signal cables, to be checked with a differential probe, will be ready next week.

- On several LV connectors at the panel side some of the shield wires got disconnected from the connector (third picture). Nick proposed, when fixing these connectors) to pre-tin the shield wires so that they could not come out one of the connector by one as they did.


- Beni presented his results from the mini-DAQ (E-log entry 125) for the wire readout. Since the TDCs work very fast one can get high statistics result in about 2 hours as shown on the plot. For the strips the readout is very slow, the fADC125 firmware for the fast readout is not ready yet and there's no zero suppression. Beni showed results (E-log entry 124) taken over the long weekend. One can see some hot channels and some missing. Lubomir looked with a scope at all the problematic channels and didn't see high rate in the hot channels; the missing channels were most likely due to problems with the 3M connector; Chris will look at this.

- Can we use the mini-DAQ as it is, without external trigger, for the final FDC test? Everybody agrees that if the fast readout is not ready we can't. Even if it is ready we still will not be able to have timing spectra (like drift times) without external trigger that may be needed the check for example if the negative HV is OK. However (Eugene) with the chambers in a vertical position it is not clear if the drift time spectra will make sense.


- Next week meeting, only, will be in B101.